OpenStudy (anonymous):

Select the factors of x2 − 10x + 25

OpenStudy (anonymous):

(x-5)(x-5)

OpenStudy (anonymous):

x2-10x+25 =(x)^2-2(x)(5)+(5)^2 =(x-5)^2